A heart dysfunction that can cause death is ventricular fibrillation. This is an uncoordinated quivering of the heart as opposed to regular beating. An electric shock to the chest can cause momentary paralysis of the heart muscle, after which the heart will sometimes start organized beating again. A defibrillator is a device that applies a strong electric shock to the chest over a time of a few milliseconds. The device contains a capacitor of several microfarads, charged to several thousand volts. Electrodes called paddles, about 8 cm across and coated with conducting paste, are held against the chest on both sides of the heart, their handles are insulated to prevent injury to the operator, who calls clear and pushes a button on one paddle to discharge the capacitor through the patients chest.

assume that the capacitor in the defibrillator is 10 uF and is charged to 5,000 volts.

a. how much charge is stored in the capacitor before it is discharged?
b. how much energy is released when the capacitor is discharged?
c. if the capacitor completely discharges in 2.5 ms, what is the average current delivered by the defibrillator?
d. what is the average power delivered?
Answer #1
a)

Q = CV = 5000*10x10-6 = 0.05 C

b)

Energy = 1/2 CV^2 = 1/2* 10x10-6 * (5000)^2 = 125 J

c)

average current = average charge / average time = 0.05/2.5x10-3 = 20A

d)

average power = VI = 5000*20 = 10^5 Watts

let me know if you dont understand